OpenStudy (anonymous):

Find the limit to 0 of.... [(e^2x)-1]/x Is it possible to do this w/out a calculator?

OpenStudy (zzr0ck3r):

do you know l`hopital rule yet?

OpenStudy (anonymous):

no... :(

OpenStudy (anonymous):

You can't have a zero in the denominater. The answer is 2

OpenStudy (zzr0ck3r):

dont just give answer, help them find answer....I dont know how to solve this wthout lhopitals rule

OpenStudy (anonymous):

@zzr0ck3r yeah, I'm trying to find a direct way to do this, without l'Hopital's

OpenStudy (anonymous):

idk, I'm just practicing for a test and this question was online...

OpenStudy (zzr0ck3r):

you nee to know how to tke a derivative to do this, you will lear that soon

OpenStudy (anonymous):

I know how to take a derivative, but how will that help me solve it?

OpenStudy (zzr0ck3r):

take the derivative of the num and the the denom. then run the limit

OpenStudy (zzr0ck3r):

you will have (2e^2x)/1 now run the limi

OpenStudy (zzr0ck3r):

limit*

OpenStudy (anonymous):

Oh, I've got one: Well we know: \[ e^x-1\sim x\\ x\to 0 \] So: \[ \lim_{x\to 0}\frac{e^{2x}-1}{x}=\lim_{u\to 0}\frac{e^{u}-1}{\frac{1}{2}u} \]Where \(u=2x\) Due to that: \[ \lim_{x\to 0}\frac{u}{\frac{1}{2}u}=\frac{1}{\frac{1}{2}}=2 \]